# Identification and estimation of mediational effects of longitudinal modified treatment policies

### v3

@@ 6 Acknowledgments
The two terms of this contrast arise from two causal models where the mediator variables are identical. Therefore, any pathway $A_{i}\rightarrow M_{j}\rightarrow\mathbf{V}\rightarrow Y$ is blocked and cannot give rise to differences in $Y$, where $\mathbf{V}$ is any set of nodes that doesn’t include a treatment node. More precisely, any node after $M_{j}$ which is not a treatment node is, for both terms of the contrast, generated from a causal model with $M_{j}$ arising from $\mathsf{d}^{1}$, and therefore there can be no differences through such a path. ∎
This means, in other words, that effects that have a mediator variable as the proximate interventional cause as the outcome are not captured in the natural direct effect.
@@
A considerable body of literature has studied semiparametric estimation of total causal effects of static and dynamic interventions using sequential regression (e.g., Robins, 2000; van der Laan and Robins, 2003; Bang and Robins, 2005; van der Laan and Rubin, 2006; van der Laan and Rose, 2011; Luedtke et al., 2017; Rotnitzky et al., 2017; van der Laan and Rose, 2018). Typically, the approach involves the estimation of both the regression functions (in this case, the prediction of $Z,M,$ and $L$ conditional on their past) and the probability of treatment. These procedures have the property that the rate of convergence of the estimator for the treatment effect is equal to the product of the rates of convergences of the outcome and treatment estimators; this allows flexible machine learning models that converge at rates slower than “parametric” $\sqrt{n}$ rate, as long as their product converges at $\sqrt{n}$ rate (e.g., each nuisance parameter can be estimated at $n^{1/4}$ rate). In particular, if either model is consistent, then the treatment effect estimator is consistent.
The requisite $n^{1/4}$-rate estimator convergence is achievable for some certain classes of flexible predictors, including LASSO (Bickel et al., 2009), highly adaptive LASSO (Benkeser and Van Der Laan, 2016), random forests (Wager and Walther, 2015), and neural networks (Chen and White, 1999) though we have not verified this fact for each estimator implemented in Super Learner. One may also consult the discussion of convergence rates for machine learning estimators in Section 3 of (Chernozhukov et al., 2018). We also note that if (correctly specified) parametric models are used, then the assumption can be easily verified. Overall, while the theoretical convergence rates of nonparametric estimators may be difficult to verify, we recommend the doubly robust procedure we have presented because it provides the best chance for appropriate uncertainty quantification while minimizing the possibility of model misspecification.
As given below, the algebraic form of the efficient influence function for our parameter is identical to that of the parameter studied in Díaz et al. (2023b), except that the definitions of the IPW functions ($\mathsf{G}^{\prime}$ and $\mathsf{G}^{\star}$) and the iterative regressions ($\mathsf{Q}$-functions) change to accommodate the LMTP generalization. The proofs and arguments for all the following claims mirror those of Díaz et al. (2023b) and are therefore omitted here. To give an expression for the efficient influence function for $\theta$, we must introduce the following functions, which can be understood intuitively as doubly robust analogs of the $\mathsf{Q}$- functions (see Rubin and van der Laan (2007)).
@@
$\mathsf{D}_{M,t}(\underaccent{\bar}{X}_{t},\underaccent{\bar}{m}_{t})=\sum_{s=t}^{\tau}\mathsf{K}^{*}_{t,s}[\prod_{k=t}^{s-1}\mathds{1}(M_{k}=m_{k})][\mathds{1}(M_{s}=m_{s})\mathsf{Q}_{M,s+1}-Q_{M,s}]+Q_{M,t}$ (S3)
@@
Proof See the proof of Theorem 2 of Díaz et al. (2023b). ∎
The algorithm used to estimate $\theta$ is given as Algorithm 1 in Díaz et al. (2023b), where in our case the IPW weights are replaced with probability ratios as described in Section 3.2. The algorithm consists of the following steps. First, we set $\hat{\mathsf{D}}_{Z,\tau+1}=Y$ and $\hat{\mathsf{D}}_{M,\tau+1}=1$. Then, beginning at $t=\tau$, we calculate $\hat{\mathsf{Q}}_{L,t}$ as the nonparametric regression prediction of $\hat{\mathsf{D}}_{Z,t+1}$, onto preceding variables. (The method is agnostic as to the choice of regression technique, but we recommend nonparametric or “data-adaptive” approaches; see Section 4.3 of the main text for the specific tools we use in our data analysis.) Then, we nonparametrically estimate the functions $\mathsf{g}$ and $\mathsf{G}$ (as in Section 3.2), and $\hat{\mathsf{D}}_{L,t}$ is calculated by plugging in all the nuisance functions estimated so far into its definition (in equation S1). Similarly, $\hat{\mathsf{Q}}_{Z,t}$ comes from a nonparametric regression of $\hat{\mathsf{D}}_{L,t}$, and $\hat{\mathsf{Q}}_{M,t}$ comes from a nonparametric regression of $\mathbbm{I}(M_{t}=m_{t})\hat{\mathsf{D}}_{M,t+1}$. Then $\hat{\mathsf{D}}_{Z,t}$ is estimated from equation S2 and $\hat{\mathsf{D}}_{M,t}$ is estimated from equation S3. This is repeated for $t=\tau-1,\tau-2,\dots,1$. Finally, the mean of $\hat{\mathsf{D}}_{Z,1}$ is used as an estimator for $\phi$, and the mean of $\hat{\mathsf{D}}_{M,1}$ is taken as an estimator of $\lambda$, which are plugged into the identifying formula of Theorem 1 to estimate $\theta$. (Note that the use of the $\mathsf{D}$-functions in this way is motivated by Lemma 1 of Díaz et al. (2023b).) The estimated variance of the efficient influence function (Theorem 1) is used to estimate the standard error. The procedure makes use of cross-fitting in all regression functions to ensure that the asymptotic linearity result above holds without restrictions on the complexity of the form of the machine learning estimators, thereby yielding valid statistical inference under mild conditions (Zheng and van der Laan, 2011; Chernozhukov et al., 2018). See Section 5 of Díaz et al. (2023b) for additional discussion of the convergence of the estimator described in that paper, which also applies to the current case.
S6 More on the structure of the observed data and intervention
@@
Although we are mainly scientifically interested in the intervention to delay intubation, we re-iterate here that part of the intervention we consider is to “prevent censoring”; this intervention applies to both treatment arms we are comparing (status quo and delay intubation). What this means practically is that outcomes are estimated for censored observations based on our estimates of the nuisance functions. While this may introduce some uncertainty relative to an uncensored data set, it does not induce bias unless the censoring at random assumption is violated. While censored individuals might be thought a priori to be, for example, more healthy (if they are discharged), our goal has been to control for enough baseline and time-varying covariates to make such conditional differences minimal. Statistically, our procedure has the effect that censored observations do not provide direct information about the effect of the intervention to delay intubation (because they do not provide outcome data), but they do inform the estimates of nuisance functions that contribute to the estimation of the effect.
@@
Often, “survival analysis” is performed by analyzing differences in hazard functions, either semi-parametrically through a proportional hazards assumption and Cox regression or through distribution-free hypothesis testing with the logrank test. Although such methods are plausible, in the context of causal inference it is more natural to define a specific parameter that quantifies the scientific effect to be estimated. In this case, the parameter is the difference in mortality rates between two different treatment regimes: one the status quo intervention, and one the intervention with delayed intubation. Thus, the form of the parameter is fundamentally no different from treatment effects in non-survival contexts except that censoring is considered. To add extra temporal dimensions to the results, one could repeat the analysis for multiple values of follow-up time or delay time, but we have chosen to estimate just one parameter for the current study.
S7 Effect modification in the AKI study
Here, we aim to identify treatment effect modifiers that can be used to predict which patients are at a higher risk of a harmful effect through AKI, and which patients are likely to benefit due to all other mechanisms. Rigorously estimating interpretable conditional average treatment effects, or treatment effects given some covariate values, is notoriously difficult, especially when continuous covariates are considered (see for example Kennedy et al. (2017) for a review of the state-of-the-art in estimation of conditional effects). For simplicity and interpretability, we can approach this by following a strategy presented in Rudolph and Díaz (2022). Specifically, we regress the influence function (see Theorem 1) for the interventional direct effect of intubation onto baseline covariates. (It is easily shown that the conditional (on covariates) expectation of the (uncentered) influence function for the ATE in standard settings is equal to the conditional average treatment effect; for the extension of this principle to the LMTP case see equation 6 of Díaz et al. (2021). To derive the influence function of the direct and indirect effects, note that the influence function for the difference of two parameters is equal to the difference of the influence functions of those parameters.)
@@
First, we note that this analysis is exploratory; some subsets may have relatively small sample sizes and we have not calculated standard errors. However, interstitial lung disease (ILD) is the most important variable for effect modification for both the indirect effect through AKI and the direct effect through other mechanisms. The negative effect modification parameter for ILD in Table S2 implies that delaying IMV in patients with ILD has a worse effect on survival (i.e., is more harmful/less beneficial) due to mechanisms other than AKI than in patients without ILD, whereas the positive effect modification parameter in Table S2 means that delaying IMV in patients with ILD has a better effect on survival (i.e., is less harmful/more beneficial) due to AKI than in patients without ILD. The former effect can be explained by the fact that delaying intubation may be riskier in patients with ILD; the latter effect (smaller in size) might be due to the fact that patients with ILD are more likely to experience mortality due to lung failure rather than AKI.
Likewise, Table S2 shows that delaying IMV in patients with cancer, chronic obstructive pulmonary disease (COPD), high lymphocyte count, costovertebral angle tenderness (CVA), coronary artery disease (CAD), high creatinine values, and those who used home oxygen, has a better effect on survival (i.e., is less harmful/more beneficial) due to AKI than in patients without those conditions. Delaying IMV in patients who have cirrhosis or who are current smokers has a worse effect on survival (i.e., is more harmful/less beneficial) due to AKI than in patients without those conditions.
The results in Table S2 show that delaying intubation in patients who are immunosuppressed, use home oxygen, or are white, has a worse effect on survival (i.e., is more harmful/less beneficial) due to mechanisms other than AKI than in patients outside those categories. Lastly, delaying intubation in patients with high creatinine values, asthma, CVA, cirrhosis, high bilirubin, and who are former smokers, has a better effect on survival (i.e., is less harmful/more beneficial) due to mechanisms other than AKI than in patients without those conditions.
@@
Table S2: Most important effect modifiers for the direct and indirect effect of invasive mechanical ventilation on survival through acute kidney failure, as measured by the slope of a simple linear regression of the efficient influence function. Variables with negative slopes indicate populations with lower effects (calculated as intervention minus baseline; positive values indicate that delaying intubation is relatively less harmful/more beneficial).
@@
(a) Effect modifiers for the effect of delaying IMV on survival through AKI.
@@
(b) Effect modifiers for the effect of delaying IMV on survival independent of AKI.
S8 Simulation study
To test the accuracy of our estimator, we undertook a simulation study described by the following data-generating distribution.
$L_{1}$ $\sim\mathcal{N}(0,1)$
$A_{1}$ $\sim\text{Bernoulli}(\frac{1}{3}+\frac{1}{3}\text{logit}^{-1}(0.5L_{1}))$
$Z_{1}$ $\sim\mathcal{N}(0.5(-L_{1}+A_{1}-0.5),1)$
$M_{1}$ $\sim\text{Bernoulli}(\frac{1}{3}+\frac{1}{3}\text{logit}^{-1}(0.5(L_{1}-A_{1}-Z_{1}+0.5)))$
$L_{2}$ $\sim\mathcal{N}(0.5(-L_{1}+A_{1}+Z_{1}-M_{1}),1)$
$A_{2}$ $\sim\text{Bernoulli}(\frac{1}{3}+\frac{1}{3}\text{logit}^{-1}(0.5(L_{1}-A_{1}-Z_{1}+M_{1}+L_{2})))$
$Z_{2}$ $\sim\mathcal{N}(0.5(-L_{1}+A_{1}+Z_{1}-M_{1}-L_{2}+A_{2}-0.5),1)$
$M_{2}$ $\sim\text{Bernoulli}(\frac{1}{3}+\frac{1}{3}\text{logit}^{-1}(0.5(L_{1}-A_{1}-Z_{1}+M_{1}+L_{2}-A_{2}-Z_{2}+0.5)))$
$Y$ $\sim\mathcal{N}(0.5(-L_{1}-A_{1}-Z_{1}-M_{1}+L_{2}+UA_{2}-Z_{2}+VM_{2}),1)$
where $U,V\in{-1,1}$ are parameters that vary across simulation scenarios. We vary the sample size $n\in(1000,5000)$. Each combination of $U,V,n$ is run for 300 replicates.
We consider two treatment policies; $\mathsf{d}$ is the status quo (according to the data-generating process), and $\mathsf{d}^{\prime}$ is the intervention that sets $A_{1}=1$ and then samples $A_{2}$ naturally, according to the data-generating process. Our estimand of interest is $\theta=\mathsf{E}[Y(\bar{A}^{\mathsf{d}^{\prime}},J(\bar{A}^{d})]$ (see main paper for an explanation of this notation). The true value of the target parameter is $\theta(U,V)=(U-1.25+V/2)/2$.
Estimation was performed using the lcmmtp package with cross-fitting ($K=3$ folds) and an ensemble of learners in the Super Learner framework (van der Laan et al., 2007) including mean, GLM, multivariate adaptive regression splines, and random forest. For each parameter combination $(n,U,V)$, 300 simulation replicates were conducted.
$n$ $U$ $V$ $\theta$ $n\cdot\text{MSE}$ Coverage
$1\cdot 10^{3}$1 -1 -1.375 3.872 0.950
$5\cdot 10^{3}$1 -1 -1.375 4.039 0.930
$1\cdot 10^{3}$1 1 -0.875 3.255 0.963
$5\cdot 10^{3}$1 1 -0.875 3.797 0.953
$1\cdot 10^{3}$ 1 -1 -0.375 4.486 0.920
$5\cdot 10^{3}$ 1 -1 -0.375 3.827 0.943
$1\cdot 10^{3}$ 1 1 0.125 3.296 0.947
$5\cdot 10^{3}$ 1 1 0.125 3.591 0.963
Table S3: Summary of simulation scenarios and results. The table presents the sample size ($n$), parameters $U$ and $V$, the true parameter value ($\theta$), the product of the sample size and mean squared error ($n\cdot\text{MSE}$), and the coverage probability for each scenario in the simulation study.
[…]
Figure S3: Coverage Probability for the parameter estimate across different sample sizes and scenarios. Lines connect points with the same $U$ and $V$ values.
As shown in the figures and table, the method demonstrates good accuracy (with steady $n\cdot\text{MSE}$ across sample sizes) and approximately nominal coverage. As a final remark, we note that we have chosen a relatively simple data-generating process, making use of linear functions, in order to facilitate the calculation of the theoretical value of $\theta$, but those functional forms are not assumed by the model or estimators, which can accommodate much more complex data-generating processes.
Code for the simulation study can be found at https://github.com/bjg345/lcmmtp_sim/.

### v4

@@ 6 Acknowledgments
The two terms of this contrast arise from two causal models where the mediator variables are identical. Therefore, any pathway $A_{i}\rightarrow M_{j}\rightarrow\mathbf{V}\rightarrow Y$ is blocked and cannot give rise to differences in $Y$, where $\mathbf{V}$ is any set of nodes that does not include a treatment node. More precisely, any node after $M_{j}$ which is not a treatment node is, for both terms of the contrast, generated from a causal model with $M_{j}$ arising from $\mathsf{d}^{1}$, and therefore there can be no differences through such a path. ∎
This means, in other words, that effects that have a mediator variable as the proximate interventional cause of the outcome are not captured in the natural direct effect.
@@
A considerable body of literature has studied semiparametric estimation of total causal effects of static and dynamic interventions using sequential regression (e.g., Robins, 2000; van der Laan and Robins, 2003; Bang and Robins, 2005; van der Laan and Rubin, 2006; van der Laan and Rose, 2011; Luedtke et al., 2017; Rotnitzky et al., 2017; van der Laan and Rose, 2018). Typically, the approach involves the estimation of both the regression functions (in this case, the prediction of $Z,M,$ and $L$ conditional on their past) and the probability of treatment. These procedures have the property that the rate of convergence of the estimator for the treatment effect is equal to the product of the rates of convergence of the outcome and treatment estimators; this allows flexible machine learning models that converge at rates slower than “parametric” $\sqrt{n}$ rate, as long as their product converges at $\sqrt{n}$ rate (e.g., each nuisance parameter can be estimated at $n^{1/4}$ rate). In particular, if either model is consistent, then the treatment effect estimator is consistent.
The requisite $n^{1/4}$-rate estimator convergence is achievable for some certain classes of flexible predictors, including LASSO (Bickel et al., 2009), highly adaptive LASSO (Benkeser and Van Der Laan, 2016), random forests (Wager and Walther, 2015), and neural networks (Chen and White, 1999), though we have not verified this fact for each estimator implemented in Super Learner. One may also consult the discussion of convergence rates for machine learning estimators in Section 3 of Chernozhukov et al. (2018). We also note that if (correctly specified) parametric models are used, then the assumption can be easily verified. Overall, while the theoretical convergence rates of nonparametric estimators may be difficult to verify, we recommend the doubly robust procedure we have presented because it provides the best chance for appropriate uncertainty quantification while minimizing the possibility of error due to model misspecification.
As given below, the algebraic form of the efficient influence function for our parameter is identical to that of the parameter studied in D´ıaz et al. (2023b), except that the definitions of the IPW functions ($\mathsf{G}^{\prime}$ and $\mathsf{G}^{\star}$) and the iterative regressions ($\mathsf{Q}$-functions) change to accommodate the LMTP generalization. The proofs and arguments for all the following claims mirror those of D´ıaz et al. (2023b) and are therefore omitted here. To give an expression for the efficient influence function for $\theta$, we must introduce the following functions, which can be understood intuitively as doubly robust analogs of the $\mathsf{Q}$- functions (see Rubin and van der Laan (2007)).
@@
$\mathsf{D}_{M,t}(\underaccent{\bar}{X}_{t},\underaccent{\bar}{m}_{t})=\sum_{s=t}^{\tau}\mathsf{K}^{*}_{t,s}[\prod_{k=t}^{s-1}\mathbb{I}(M_{k}=m_{k})][\mathbb{I}(M_{s}=m_{s})\mathsf{Q}_{M,s+1}-\mathsf{Q}_{M,s}]+\mathsf{Q}_{M,t}$ (S3)
@@
Proof See the proof of Theorem 2 of D´ıaz et al. (2023b). ∎
The algorithm used to estimate $\theta$ is given as Algorithm 1 in D´ıaz et al. (2023b), where in our case the IPW weights are replaced with probability ratios as described in Section 3.2. The algorithm consists of the following steps. First, we set $\hat{\mathsf{D}}_{Z,\tau+1}=Y$ and $\hat{\mathsf{D}}_{M,\tau+1}=1$. Then, beginning at $t=\tau$, we calculate $\hat{\mathsf{Q}}_{L,t}$ as the nonparametric regression prediction of $\hat{\mathsf{D}}_{Z,t+1}$, onto preceding variables. (The method is agnostic as to the choice of regression technique, but we recommend nonparametric or “data-adaptive” approaches; see Section 4.3 of the main text for the specific tools we use in our data analysis.) Then, we nonparametrically estimate the functions $\mathsf{g}$ and $\mathsf{G}$ (as in Section 3.2), and $\hat{\mathsf{D}}_{L,t}$ is calculated by plugging in all the nuisance functions estimated so far into its definition (in equation S1). Similarly, $\hat{\mathsf{Q}}_{Z,t}$ comes from a nonparametric regression of $\hat{\mathsf{D}}_{L,t}$, and $\hat{\mathsf{Q}}_{M,t}$ comes from a nonparametric regression of $\mathbb{I}(M_{t}=m_{t})\hat{\mathsf{D}}_{M,t+1}$. Then $\hat{\mathsf{D}}_{Z,t}$ is estimated from equation S2 and $\hat{\mathsf{D}}_{M,t}$ is estimated from equation S3. This is repeated for $t=\tau-1,\tau-2,\dots,1$. Finally, the mean of $\hat{\mathsf{D}}_{Z,1}$ is used as an estimator for $\varphi$, and the mean of $\hat{\mathsf{D}}_{M,1}$ is taken as an estimator of $\lambda$, which are plugged into the identifying formula of Theorem 1 to estimate $\theta$. (Note that the use of the $\mathsf{D}$-functions in this way is motivated by Lemma 1 of D´ıaz et al. (2023b).) The estimated variance of the efficient influence function (Theorem 1) is used to estimate the standard error. The procedure makes use of cross-fitting in all regression functions to ensure that the asymptotic linearity result above holds without restrictions on the complexity of the form of the machine learning estimators, thereby yielding valid statistical inference under mild conditions (Zheng and van der Laan, 2011; Chernozhukov et al., 2018). See Section 5 of D´ıaz et al. (2023b) for additional discussion of the convergence of the estimator described in that paper, which also applies to the current case.
Appendix S6 More on the structure of the observed data and intervention
@@
Although we are mainly scientifically interested in the intervention to prevent intubation, we reiterate here that part of the intervention we consider is to “prevent censoring”; this intervention applies to both treatment arms we are comparing (status quo and prevent intubation). What this means practically is that outcomes are estimated for censored observations based on our estimates of the nuisance functions. While this may introduce some uncertainty relative to an uncensored data set, it does not induce bias unless the censoring at random assumption is violated. While censored individuals might be thought a priori to be, for example, healthier (if they are discharged), our goal has been to control for enough baseline and time-varying covariates to make such conditional differences minimal. Statistically, our procedure has the effect that censored observations do not provide direct information about the effect of the intervention to prevent intubation (because they do not provide outcome data), but they do inform the estimates of nuisance functions that contribute to the estimation of the effect.
@@
Often, “survival analysis” is performed by analyzing differences in hazard functions, either semi-parametrically through a proportional hazards assumption and Cox regression or through distribution-free hypothesis testing with the logrank test. Although such methods are plausible, in the context of causal inference it is more natural to define a specific parameter that quantifies the scientific effect to be estimated. In this case, the parameter is the difference in mortality rates between two different treatment regimes: one the status quo intervention, and one the intervention with prevented intubation. Thus, the form of the parameter is fundamentally no different from treatment effects in non-survival contexts except that censoring is considered. To add extra temporal dimensions to the results, one could repeat the analysis for multiple values of follow-up time or consider alternative forms of the intervention.
Appendix S7 Effect modification in the AKI study
Here, we aim to identify treatment effect modifiers that can be used to predict which patients are at a higher risk of a harmful effect through AKI, and which patients are likely to benefit due to all other mechanisms. Rigorously estimating interpretable conditional average treatment effects, or treatment effects given some covariate values, is notoriously difficult, especially when continuous covariates are considered (see for example Kennedy et al. (2017) for a review of the state-of-the-art in estimation of conditional effects). For simplicity and interpretability, we can approach this by following a strategy presented in Rudolph and D´ıaz (2022). Specifically, we regress the influence function (see Theorem 1) for the interventional direct effect of intubation onto baseline covariates. (It is easily shown that the conditional (on covariates) expectation of the (uncentered) influence function for the ATE in standard settings is equal to the conditional average treatment effect; for the extension of this principle to the LMTP case see equation 6 of D´ıaz et al. (2021). To derive the influence function of the direct and indirect effects, note that the influence function for the difference of two parameters is equal to the difference of the influence functions of those parameters.)
@@
First, we note that this analysis is exploratory; some subsets may have relatively small sample sizes and we have not calculated standard errors. However, interstitial lung disease (ILD) is the most important variable for effect modification for both the indirect effect through AKI and the direct effect through other mechanisms. The negative effect modification parameter for ILD in Table S2 implies that preventing IMV in patients with ILD has a worse effect on survival (i.e., is more harmful/less beneficial) due to mechanisms other than AKI than in patients without ILD, whereas the positive effect modification parameter in Table S2 means that preventing IMV in patients with ILD has a better effect on survival (i.e., is less harmful/more beneficial) due to AKI than in patients without ILD. The former effect can be explained by the fact that preventing intubation may be riskier in patients with ILD; the latter effect (smaller in size) might be due to the fact that patients with ILD are more likely to experience mortality due to lung failure rather than AKI.
Likewise, Table S2 shows that preventing IMV in patients with cancer, chronic obstructive pulmonary disease (COPD), high lymphocyte count, costovertebral angle tenderness (CVA), coronary artery disease (CAD), high creatinine values, and those who used home oxygen, has a better effect on survival (i.e., is less harmful/more beneficial) due to AKI than in patients without those conditions. Preventing IMV in patients who have cirrhosis or who are current smokers has a worse effect on survival (i.e., is more harmful/less beneficial) due to AKI than in patients without those conditions.
The results in Table S2 show that preventing intubation in patients who are immunosuppressed, use home oxygen, or are white, has a worse effect on survival (i.e., is more harmful/less beneficial) due to mechanisms other than AKI than in patients outside those categories. Lastly, preventing intubation in patients with high creatinine values, asthma, CVA, cirrhosis, high bilirubin, and who are former smokers, has a better effect on survival (i.e., is less harmful/more beneficial) due to mechanisms other than AKI than in patients without those conditions.
@@
Table S2: Most important effect modifiers for the direct and indirect effect of invasive mechanical ventilation on survival through acute kidney injury, as measured by the slope of a simple linear regression of the efficient influence function. Variables with negative slopes indicate populations with lower effects (calculated as intervention minus baseline; positive values indicate that preventing intubation is relatively less harmful/more beneficial).
@@
(a) Effect modifiers for the effect of preventing IMV on survival through AKI.
@@
(b) Effect modifiers for the effect of preventing IMV on survival independent of AKI.
Appendix S8 Simulation study
To test the accuracy of our estimator, we performed a brief Monte Carlo simulation study. For $t=1,2,3$ and sample sizes $n\in\{1000,5000\}$, we drew $J=300$ datasets from the following data-generating mechanism, where $Y_{1}=1$.
$L_{t}$ $\sim\text{Bernoulli}(0.5)$
$A_{t}$ $\sim\text{Bernoulli}(0.4+0.1\,L_{t})$
$M_{t}$ $\sim\text{Bernoulli}(0.6-0.1\,A_{t}+0.1\,L_{t})$
$Y_{t+1}$ $\sim\text{Bernoulli}(Y_{t}\,0+(1-Y_{t})\,(0.3+0.05\,M_{t}+0.2\,A_{t}-0.001\,L_{t}))$
Note that $Y_{t}$ is degenerate and analogous to a survival outcome. For each dataset, we estimated $\theta(\mathsf{d}^{\prime},\mathsf{d})=\mathsf{E}[Y(\bar{A}^{\mathsf{d}^{\prime}},J(\bar{A}^{\mathsf{d}}))]$ (see main paper for an explanation of this notation) under three parameterizations of $(\mathsf{d}^{\prime},\mathsf{d})$: $\hat{\theta}(1,1)$, $\hat{\theta}(1,0)$, $\hat{\theta}(0,0)$. We then estimated the randomized total, indirect, and direct effects as $\hat{\theta}(1,1)-\hat{\theta}(0,0)$, $\hat{\theta}(1,1)-\hat{\theta}(1,0)$, and $\hat{\theta}(1,0)-\hat{\theta}(0,0)$ respectively. The true values were estimated by drawing the counterfactuals $Y(\bar{A}^{\mathsf{d}^{\prime}},J(\bar{A}^{\mathsf{d}}))$ from a super-population. Estimation was performed using the lcmmtp package in R; the estimator used cross-fitting ($K=5$ folds) and a generalized linear model with all possible 3-way interactions for fitting nuisance parameters. Code for the simulation is available on GitHub at https://github.com/nt-williams/lcmmtp-simulation.
$n$ Effect $\theta$ Abs. bias $n\cdot\text{MSE}$ Coverage
1000 Total -0.219 0.024 551.388 0.970
Indirect 0.004 0.008 16.585 0.977
Direct -0.223 0.032 503.256 0.967
5000 Total -0.219 0.002 4.815 0.937
Indirect 0.004 0.000 0.392 0.923
Direct -0.223 0.002 3.726 0.947
Table S3: Summary of simulation results. The table presents the sample size ($n$), the true parameter value ($\theta$), absolute bias, the product of the sample size and mean squared error ($n\cdot\text{MSE}$), and the coverage probability for the total, indirect, and direct effect in the simulation study.
Appendix S9 Event counts in illustrative study
Table S4: Distribution of Oxygenation Levels (Days 1–4)
Day 1 2 3 4
No oxygen 1267 470 188 102
Non-invasive 1921 2546 2477 2184
Mechanical vent 176 249 304 349
Died or discharged 0 99 395 729
Table S5: Event Counts by Day (AKI, Mortality, Discharged)
Day 1 2 3 4
AKI 474 588 592 622
Mortality 24 72 59 72
Discharged 75 323 670 1003